# Timeseria: an object-oriented time series processing library

### v2

@@ 2.2 Software functionalities
As already introduced, Timeseria provides functionalities for loading, storing, manipulating and plotting data, as well as fitting, evaluating and applying models. In the following, we provide a breakdown of the main ones, together with their respective methods or classes.
@@ 2.2.1 Loading and storing data
Timeseria supports loading and storing data directly from the TimeSeries class. The from_csv() method allows to load CSV files with automatic encoding and format detection, and supports many parameters should the auto-detection fail. Field and newline separators, timestamp formats, timestamp columns and which data columns to read can all be set, as well as many others. Time series can be saved either as standard CSV files using the to_csv() method, or as dedicated format with the save() method, that includes some metadata and that fully supports data indexes. This is particularly useful for loading time series within Timeseria itself later on, which can be done using the load() method. Lastly, Pandas DataFrames are supported as well, using the to_df() and from_df() methods.
@@ 2.2.2 Manipulating and inspecting data
Data can be manipulated using the transformations and the operations, which can be accessed directly as methods of the TimeSeries class. Besides the two main resample() and aggregate() transformations, Timeseria provides a number of built-in operations which include: min(), max(), avg(), sum(), derivative(), integral(), diff(), csum(), mavg(), normalize(), offset() and rescale() to operate on the time series values, plus filter() to filter by label, slice() to cut a portion of a time series and merge() to merge two or more time series.
The TimeSeries class also supports the square brackets notation for quickly accessing its elements either by position or timestamp (e.g. timeseries[2], timeseries[dt(2015,10,25,6,19,0)]), or for filtering on a specific data label if using key-value data (e.g. timeseries[’temperature’]). Slicing is also possible, either by position or timestamp, using the standard Python slicing notation (e.g timeseries[dt(2015,1,1,0,0,0):dt(2016,1,1,0,0,0)]).
Lastly, plotting a time series can be achieved by simply invoking the plot() method of the TimeSeries class, which generates interactive plots using a JavaScript-based engine. Such plots fully support Timeseria data structures, including the data indexes, and can be saved both as images and as self-contained HTML pages in order to be easily shared. Above a given amount of data, the plotting engine automatically performs an aggregation step, in order to speed up the plotting process and to prevent overwhelming interactive analysis environments (and web browsers) with too much data. In this case, the plot is rendered as a line chart overlying an area chart which represents the dispersion of the original data, thus allowing to retain the visual information about peaks that would be otherwise smoothed.
@@ 2.2.3 Fitting, evaluating and applying models
Timeseria provides a set of built-in models to be intended more for demonstrative and/or simple tasks rather than for real applications. These include the PeriodicAverageForecaster, the PeriodicAverageReconstructor and the PeriodicAverageAnomalyDetector, which are all based on computing periodic averages over historical data; the AARIMAForecaster and the AARIMAReconstructor, based on automatic autoregressive integrated moving averages; and the LSTMForecaster and the LSTMAnomalyDetector, which make use of an LSTM neural network.
The true goal of Timeseria is instead to provide a framework where to easily plug-in custom modeling logic by extending the base classes. As of today, such classes include the Forecaster, the Reconstructor, and the ModelBasedAnomalyDetector. When extending base classes, a set of methods and decorators is automatically inherited, so that only the model-specific logic needs to be coded. For example, the base Forecaster class provides the apply(), evaluate() and cross_validate() methods, together with two decorators: @Forecaster.fit_method and @Forecaster.predict_method, to be used to decorate the custom fit() and predict() methods. All common functionalities, including the save() and load() methods as well as the data consistency checks, are thus automatically handled.
@@ 3 Illustrative examples
The following examples can be reproduced in a clean Timeseria installation, either in a Jupyter environment or with a standard Python script. In this case, plots must be generated as images and saved to file, adding the following extra arguments to each plot() call: image=True, save_to="plot.png".
3.1 Load some data
Load an air temperature dataset, sampled at about ten minutes intervals, as a TimeSeries object. Timestamps are assumed on UTC as not otherwise specified. The time series is then printed and plotted.
@@ 3.2 Resample to hourly data
The time series loaded in the previous example is resampled at one hour sampling interval. Gaps are filled by linear interpolation and the data_loss index is added. Then, the time series in printed and plotted.
@@
3.3 Three-days forecast
A forecasting model based on a LSTM neural network is first cross-validated, then fitted and applied with a forecasting horizon of 72 (hourly) steps, for a total of three days. Data marked as missing is excluded from both the training and the evaluation of the model. Results are then plotted.
@@ 3.3 Three-days forecast
forecaster = LSTMForecaster(window=24, neurons=256,
@@
print(forecaster.cross_validate(timeseries, rounds=3,
    evaluate_error_metrics=["RMSE", "MAPE"]))
@@
{’humidity[RH]_RMSE_avg’: 1.168,
 ’humidity[RH]_RMSE_stdev’: 0.207,
 ’humidity[RH]_MAPE_avg’: 0.0196,
 ’humidity[RH]_MAPE_stdev’: 0.004,
 ’temperature[C]_RMSE_avg’: 0.362,
 ’temperature[C]_RMSE_stdev’: 0.039,
 ’temperature[C]_MAPE_avg’: 0.011,
 ’temperature[C]_MAPE_stdev’: 0.002}
@@ 3.4 Perform anomaly detection
This example fits and apply a simple anomaly detection model based on periodic averages. By default the anomaly detectors assumes to work in unsupervised mode, using some “sane defaults”. Data marked as missing is excluded from the anomaly detection process, in order to prevent false positives. The time series is then plotted together with the anomaly data index, which allow for visual inspection: the bigger anomaly is detected around the 25th of January, when there is an unexpected drop with respect to the “normal” behaviour.
@@
[INFO] Model(s) evaluated, now computing the error_metric distribution(s)...
@@ 3.5 Aggregate to daily data
This example aggregates the hourly time series in daily data, but first it changes the time zone to the right one for this data set. This is important for getting the daily aggregates right and properly handling the daylight saving time change in March. In the aggregation process, the minimum and maximum operations are also computed, besides the (default) average one. The time series is then plotted.
timeseries.change_tz("Europe/Rome")
timeseries.aggregate("1D", operations=["min","max","avg"]).plot()
@@ 4 Impact
We believe that Timeseria can impact many fields, and let researchers and professionals focus more on their core tasks and less on repetitive, low added-value ones, such as handling data consistency, time and calendar arithmetic, edge cases, model evaluation, inspection of the results, and more.
We also believe that a structured approach as proposed in this work can impact several aspects of time series analysis overall, as improving reproducibility, reducing false positives and negatives, preventing misleading conclusions, and in general to keep analysis process under control with the ultimate goal of fostering new research and applications.
@@ 5 Conclusions
We presented Timeseria, an object-oriented time series processing library implemented in Python. Our goal was to make it easier to manipulate time series data and to build statistical and machine learning models on top of it.
We tackled the various issues that arise in performing such tasks at their roots, defining a novel representation for time series data based on well defined logical units (objects), which we then used to build the library upon.
Thanks to this approach, we were able to handle most of the issues that arise in processing time series data by design, with particular attention to practical challenges and real-world scenarios.
Early adoption of Timeseria proved it to be a valid tool for everyday work, preventing both naive (yet hard to debug) and more substantial errors, while speeding up all the various activities.

### v3

@@ 3.2 Software functionalities
As previously mentioned, Timeseria provides functionalities for loading, storing, manipulating, and plotting time series data, as well for fitting, evaluating and applying models. In the following, we provide an overview of the main ones, along with their respective methods and classes.
@@ 3.2.1 Loading and storing data
Timeseria supports loading and storing data directly from the TimeSeries class. The from_csv() method allows to load CSV (Comma-Separated Values) files with automatic encoding and format detection, and supports many parameters if auto-detection fails. For example, field and newline separators, column mapping, and timestamp formats can all be set, as well as many others. Time series can be saved either as standard CSV files using the to_csv() method, or using a dedicated format with the save() method, that includes some metadata and that fully supports the data indexes. Such format can then be loaded using the load() method, and is particularly convenient for storing and loading time series within Timeseria. Lastly, Pandas DataFrames are also supported, using the to_df() and from_df() methods.
@@ 3.2.2 Manipulating and inspecting data
Data can be manipulated using the transformations and the operations, which can be accessed directly as methods of the TimeSeries class. Besides the two main resample() and aggregate() transformations, Timeseria provides a number of built-in operations which include: min(), max(), avg(), sum(), derivative(), integral(), diff(), csum() (for the cumulative sum), mavg() (for the moving average), normalize(), offset() and rescale(), which all operate on the data values, plus the filter() operation to filter by label, the slice() operation to cut a portion of a time series and the merge() operation to merge two or more time series.
The TimeSeries class also supports the square brackets notation for quickly accessing its elements either by position or timestamp (e.g. timeseries[2], timeseries[datetime(2015,10,25,6,19,0,tzinfo=utc)]), or filtering on a specific data label if using key-value data (e.g. timeseries["temperature"]). Also the standard standard Python slicing notation is supported, either by position or timestamp, (e.g timeseries[2:4]).
Lastly, plotting a time series can be achieved by simply invoking the plot() method of the TimeSeries class, which generates interactive plots using a hybrid Python-JavaScript engine based on Dygraphs[24]. Such plots fully support Timeseria data structures, including the data indexes, and can be saved both as images and self-contained HTML pages. Above a given amount of data, the plotting engine automatically performs an aggregation step, in order to speed up the plotting process and to prevent overwhelming interactive analysis environments (and Web browsers) with too much data. In this case, the plot is rendered as a line chart overlying an area chart which represents the dispersion of the original data, thus allowing to retain visual information about peaks that would otherwise get smoothed.
@@ 3.2.3 Fitting, evaluating and applying models
Timeseria provides a set of built-in models to be intended primarily for demonstrative purposes and simple tasks, rather than for real-world applications. Such models include: the PeriodicAverageForecaster, the PeriodicAverageReconstructor and the PeriodicAverageAnomalyDetector, which are all based on computing periodic averages over historical data; the AARIMAForecaster and the AARIMAReconstructor, based on Automatic Auto-Regressive Integrated Moving Averages; and the LSTMForecaster and the LSTMAnomalyDetector, which make use of a LSTM (Long-Short Term Memory) neural network.
The true goal of Timeseria is instead to provide a framework where to easily plug-in custom modeling logic by extending the base classes. As of today, such classes include the Forecaster, the Reconstructor, and the ModelBasedAnomalyDetector. When extending base classes, a set of methods and decorators is automatically inherited, so that only model-specific logic needs to be coded. For example, the base Forecaster class provides the apply(), evaluate() and cross_validate() methods, together with two decorators: @Forecaster.fit_method and @Forecaster.predict_method, to be used to decorate the custom fit() and predict(). All common functionalities, including the save() and load() methods, as well as the data consistency checks, are then automatically inherited.
@@ 4 Illustrative examples
The following examples can be reproduced with a clean Timeseria installation either in a Jupyter environment or with a standard Python script. In this case, plots must be generated as images and saved to file, adding the following extra arguments to each plot() call: image=True, save_to="plot.png".
The dataset used in these illustrative examples is an indoor air temperature and humidity dataset, sampled at about ten minutes intervals, which except from its size presents some of the most common challenges in time series processing. These include: a variable sampling rate, several data losses, and the dependency on “human” time trough the time zone and the daylight saving time.
4.1 Load and inspect the data
The data is loaded as the TimeSeries object, and timestamps are assumed on UTC, since not otherwise specified. The time series is then printed to quickly inspect it.
@@ 4.1 Load and inspect the data
The average values of temperature and humidity can be obtained using the avg() operation:
print(timeseries.avg())
{’humidity[RH]’: 43.870, ’temperature[C]’: 22.488}
The time series can be plotted with the built-in plotting engine, which will automatically aggregate it by a factor of ten to speed up the process:
timeseries.plot()
@@ 4.2 Resample to hourly data
The time series is resampled with a sampling interval of one hour. Gaps are filled by linear interpolation and the data_loss index is added. Then, the time series in printed and plotted.
@@
4.3 Three-day forecast
First of all, the time zone of the time series is changed to the right one for this dataset. This is important because of the strong correlation of the data with the time of the day: leaving the timestamps on UTC would not allow the model to learn such dependency accurately, given that the daylight saving time switch in March would introduce an offset of one hour.
timeseries.change_tz("Europe/Rome")
A forecasting model based on a LSTM neural network is now instantiated, cross-validated, and fitted. Then, it is applied with a forecasting horizon of 72 (hourly) steps, for a total of three days, and the results plotted. Data marked as missing is excluded from both the training and the evaluation of the model.
@@ 4.3 Three-day forecast
LSTM_forecaster = LSTMForecaster(window=24, neurons=256, \
@@
cross_validation = LSTM_forecaster.cross_validate(timeseries, \
    rounds=3, evaluate_error_metrics=["RMSE", "MAPE"]))
print(cross_validation)
@@
{’humidity[RH]_RMSE_avg’: 1.1284,
 ’humidity[RH]_RMSE_stdev’: 0.1547,
 ’humidity[RH]_MAPE_avg’: 0.01815,
 ’humidity[RH]_MAPE_stdev’: 0.0013,
 ’temperature[C]_RMSE_avg’: 0.3584,
 ’temperature[C]_RMSE_stdev’: 0.0152,
 ’temperature[C]_MAPE_avg’: 0.0118,
 ’temperature[C]_MAPE_stdev’: 0.0010}
@@ 4.4 Perform anomaly detection
This example fits and apply a simple anomaly detection model based on periodic averages. By default the anomaly detectors assumes to work in unsupervised mode, using some “sane defaults”. Data marked as missing is excluded from the anomaly detection process, in order to prevent false positives. The resulting time series is then plotted together with the anomaly data index, which allows for visual inspection: the bigger anomaly is detected around the 25th of January, when there is an unexpected drop in both temperature and humidity with respect to the “normal” behavior.
@@
[INFO] Model(s) evaluated, now computing the error distribution(s)...
@@ 4.5 Aggregate to daily data
The time series is aggregated in daily data, correctly handling the daylight saving time change in March. In the aggregation process, the minimum and maximum operations are also computed, besides the (default) average one. The time series is then plotted.
timeseries = timeseries.aggregate("1D", \
    operations=["min","max","avg"])
timeseries.plot()
@@
4.6 Using custom models
The examples above made use of Timeseria’s built-in models. However, as previously mentioned, the main goal of Timeseria is to let users plug-in their own modeling logic with minimal effort. This example thus demonstrated how to define a custom forecaster, by just implementing and decorating the fit() and predict() methods. The forecaster has a window of two elements, and a trivial logic which uses the average values computed during the fit phase averaged with the window elements to compute the prediction:
from timeseria.models import Forecaster
class MyForecaster(Forecaster):
    window = 2
    @Forecaster.fit_method
    def fit(self, series, verbose=False):
        self.data["avg"] = series.avg()
    @Forecaster.predict_method
    def predict(self, series, steps=1):
        if steps > 1:
            raise NotImplementedError()
        prediction = {}
        for label in self.data["avg"].keys():
            prediction[label] = (series[-1].data[label] + \
                                 series[-2].data[label] + \
                                 self.data["avg"][label]) / 3
        return prediction
Such forecaster inherited the apply(), evaluate(), cross_validate(), save(), and load() methods, which can all be used as already introduced. In a similar way, also custom model-based anomaly detectors can be defined, but they just need the model class to be set, as it follows:
from timeseria.models import ModelBasedAnomalyDetector
class MyModelBasedAnomalyDetector(ModelBasedAnomalyDetector):
    model_class = MyForecaster
The anomaly the detector inherited all the main functionalities from the base class, as the fit(), apply(), save(), and load() methods, and can be used as previously illustrated, with no further effort.
@@ 5 Impact
We believe that Timeseria can impact many fields, and let researchers and professionals focus more on their core tasks and less on repetitive, low added-value ones, such as handling data consistency, time and calendar arithmetic, edge cases, model evaluation, inspection of the results, and more. We also believe that a structured approach as proposed in this work can impact several aspects of time series analysis overall, as improving reproducibility, reducing false positives and negatives, preventing misleading conclusions, and in general to keep analysis process under control with the ultimate goal of fostering new research and applications.
@@ 6 Conclusions
We presented Timeseria, an object-oriented time series processing library implemented in Python. Our goal was to make it easier to manipulate time series data and to build statistical and machine learning models on top of it. We tackled the various issues that arise in performing such tasks at their roots, defining a novel representation for time series data based on well defined logical units (objects), which we then used to build the library upon. Thanks to this approach, we were able to handle most of the issues that arise in processing time series data by design, with particular attention to practical challenges and real-world scenarios.
Early adoption of Timeseria proved it to be a valid tool for everyday work, preventing both naive (yet hard to debug) and more substantial errors, while speeding up all the various activities. Future work include adding support for probabilistic data structures and models (under development), adding more built-in models, and improving the performance with Cython, Numba, PyPy or similar solutions.